Let "a(x) =" "x^3 - 2x - 4"


Which of the following is true?

  1. "a(x))" is not a polynomial
  2. The remainder when "a(x)" is divided by "x + 2" is 0
  3. "a(x)" is divisible by "x - 2"
  4. "a(-1) = 3"
  5. None of the above
1
Expert's answer
2021-07-02T10:53:32-0400

"a(x))" is not a polynomial

false because it is a polynomial.


2:

The remainder when a(x) is divided by x+2 is 0

Check

"\frac{x^3-2x-4}{x+2}"

Find the next term of the quotient and the remainder

"x^2+\frac{-2x^2-2x-4}{x+2}"


"\frac{x^3-2x-4}{x+2}"

Find the next term of the quotient and the remainder m

"\frac{x^3-2x-4}{x+2}"

Find the next term of the quotient and the remainder

"x^2-2x+\frac{2x-4}{x+2}"

Find the next term of the quotient and the remainder

"x^2-2x+2+\frac{-8}{x+2}"

Simplify the sign in the fraction

"x^2-2x+2-\frac{8}{x+2}"


False because the remainder is -8.


3: 

a(x) is divisible by x-2

Check

Use polynomial division for x

"\frac{x^3-2x-4)}{x-2}"

Find the next term of the quotient and the remainder

"x^2+\frac{-2x^2-2x-4}{x-2}"

Find the next term of the quotient and the remainder

"x^2+2x+\frac{2x-4}{x-2}"

Find the next term of the quotient and the remainder

"x^2+2x+2"


True because a(x) divide by 2 is "x^2+2x+2"


4:

"a(-1)=3\\check\\a(-1)=(-1)^3-2(-1)-4=-1+2-4=-5+2=-3\\False"